Jeremy purchased a machine for $21,000 for his company. He paid 5.00% of this amount as a down payment and financed the rest at 6.92% compounded monthly. He paid $1,125 at the end of every month to settle the loan.

a. What was the principal portion of payment number 4?


b. What was the interest portion of payment number 4?
1
Expert's answer
2020-05-21T13:07:44-0400

Jeremy’s debt for the car is initially $ 19950. (0.95*21000)

6.92% at year, 0,576% at months.

19950 $ debt. for first months

19950*0.00576=114,9 $ persent portion(pp)

Interest portion(Ip) 1125-114,9=1010,1$

19950-1010,1=18939,9$ debt fod second months:

Pp:18939*0.00576=109,1$

Ip:1125-109,1=1016,9$

18939,9-1016,9=17923$ debt for third month.

Pp:17923*0.00576=103,2$

Ip:1125-103,2=1021,8$

17923-1021,8=16901,2 debt for fourth:

Pp:16901,2*0.00576=97.4$

Ip:1125-97,4=1027,6$